ELECTRONIC SPECTROSCOPY OF CARBON CHAINS AND RELEVANCE TO ASTROPHYSICS

\underline{JOHN P. MAIER}, Institute of Physical Chemistry. University of Basel, Klingelbergstr. 80, CH-4056 Basel, Switzerland.

The electronic spectra of a variety of carbon chanis, C_{n}, C_{n}H, HC_{n}H and their ions C_{n}^{−}, HC_{n}H^{+} have been identified in neon matrices between 200 and 2500 nm. This has been achieved by mass-selected codeposition of ions with excess of neon to form a matrix at 5 K. The understanding of the electronic transitions of the carbon chains point out which type and sizes are relevant for astrophysical consideration, such as in relation to the diffuse interstellar bands. Infrared bands of some such mass-selected carbon species are also detectable. The technique has also been used to elucidate the formation of carbon chains in argon and neon matrices by diffusive coagulation on annealing.